A leather coin purse has the shape of a triangle. Two sides are equal in length and the third side is 3 centimeters shorter than one-half times the length of the equal sides. The perimeter is 28.5 centimeters. Find the lengths of the sides.
Answer by mananth(16946)

A leather coin purse has the shape of a triangle. Two sides are equal in length and the third side is 3 centimeters shorter than one-half times the length of the equal sides. The perimeter is 28.5 centimeters. Find the lengths of the sides.
...
let the equal sides be x
third side = x/2 -3 cm
x+x + 0.5x-3=28.5
2.5x=28.5+3
2.5x=31.5
/2.5
x= 12.6 equal side
third side = 0.5x-3
0.5 * 12.6 -3
=3.3 third side
...
m.ananth@hotmail.ca